# Deciphering the Milky Way disc formation time encrypted in the chrono-kinematics of the bar

### v1

@@ 6.1.1 The very metal-poor (VMP) disc
Sestito et al. (2019, 2020) found that a large fraction of VMP stars with metallicity $\mathrm{[M/H]}\lesssim-2.5$ have disc-like kinematics. Several possible scenarios for the existence of VMP stars with disc-like orbits are discussed in Sestito et al. (2020). Of these, albeit the least likely and clearly disfavoured by the authors, is that in which the Milky Way disc was already in place during the very metal-poor era. Di Matteo et al. (2020) argued that stars with disc-like kinematics can be found ubiquitously across the entire metallicity range reaching $\mathrm{[M/H]}\sim-6$, thus favouring a scenario of disc formation at very low metallicities. Note that, according to numerical simulations, the MW managed to form a coherently rotating, dominant stellar disc one of the earliest amongst the galaxies of similar mass (see e.g. Semenov et al., 2024; Dillamore et al., 2024). However, given the mass and metallicity evolution of the Galaxy, its ancient disc stars were not metal-poor.
@@
Figure 10: The effects of the age uncertainty in determining the disc formation time. The left sets of panels are for Au18, and right sets are for Au26. In each sets of figures, the top row focuses on stars on the disc, which $125<R<$ 12 and $|z|<5$ kpc, and the $v_{R}$ vs. $v_{\phi}$ plane is used to distinguish disc and halo population; the bottom row focuses on inner galaxy stars with $R<5$ kpc, and the radial velocity map is used to show the disc star appearance. Left column: the age histograms for the old stars selected from uncertainty scattered age. The distribution of the scattered age is in red, and the true age is in black. Middle column: the kinematics of the true old stars born before the spin-up. Right column: the kinematics of the old stars selected with age after scattering with $10\%$ uncertainty. The aqua contours are for highlighting the distribution of all stars in the $v_{R}$ vs. $v_{\phi}$ space, and the black contours are used to illustrate the distribution of metal-rich stars.
@@ 6.1.2 Age of the disc
Queiroz et al. (2023) applied a t-SNE map to the chemical abundances measured by APOGEE and used a clustering algorithm to select a population chemically similar to the thick disc. The isochrone ages of the thick disc stars are generally old (their distribution peaks at $\mathrm{age}\gtrsim 10.9$ Gyr). Although their results are not in direct contradiction with our constraints on the disc formation time, we discuss and demonstrate in Section 6.2 below that age uncertainty can produce an overestimation of the disc age measurements. Similarly, Miglio et al. (2021) found an age of $\sim 11$ Gyr after deconvolving the age uncertainty for the chemically-defined thick disc populations with ages from asteroseismological measurements using Kepler data. Gallart et al. (2024) designed a colour-magnitude diagram(CMD)-fitting technique to reveal the star formation history of thin disc stars in the solar neighbourhood ($d\lesssim 100$ pc). They found that the star formation of the disc started around $\sim 11-10.5$ Gyr ago. Belokurov & Kravtsov (2024) separated the globular clusters into in-situ and accreted with an energy-angular momentum cut. The in-situ globular clusters spun up, i.e. increased dramatically their bulk azimuthal velocity at $-1.3\lesssim\mathrm{[Fe/H]}\lesssim-1$ around $11.7-12.7$ Gyr ago, which is also in broad agreement with the result in this work.
Nepal et al. (2024) used the StarHorse Bayesian isochrone ages (Queiroz et al., 2018) with stellar parameters from the hybrid-CNN analysis of the Gaia RVS sample (Guiglion et al., 2024) and found a kinematically cold, thin disc-like population emerging within 1 Gyr after the Big Bang. Their old, thin disc-like population covers a large range in metallicity from metal-poor to super-metal-rich and is particularly conspicuous at super-solar metallicity. It is difficult to reach super-metal-rich metallicities in the Galactic disc region within 1 Gyr after the Big Bang. It could be therefore that some (or many) of these old disc stars have come from the inner Galaxy, and thus are related to effects of the bar (Dillamore et al., 2023; Yuan et al., 2023; Li et al., 2023).
@@ 6.2 Effects from the age uncertainty
Scatter due to the age measurement uncertainty causes younger stars to be confused for apparently old stars. As the early Galaxy built its stellar mass up and the star formation rate increased, significantly more stars formed during and after the spin-up of the Milky Way than before. Therefore, the genuinely old stars are a relative minority compared to the subsequent stellar generations. These younger stars can produce noticeable contamination in a given bin of measured old age if the age uncertainty is non-negligible. We demonstrate the effects of such contamination using model galaxies Au18 and Au26 introduced earlier, as shown in Fig. 10.
@@
To illustrate how this contamination can cause overestimation of the disc age, we compare the kinematics of the genuinely old stars to that of the age-scattered old stars. In the disc region, we use the $v_{R}$ vs. $v_{\phi}$ map to distinguish the halo and the disc population. Focusing on the genuinely old population we see that before the galaxy’s spin-up, no rotation-supported component is found, as expected. However, using the same age bin but now applied to stars with scattered ages, a fast-rotating, disc-like population appears, which can be seen from the aqua contours of the number distribution in the $v_{R}$ vs. $v_{\phi}$ map. This apparently old disc-like component is solely due to the contamination by the younger stars. The disc population is more conspicuous when we further select metal-rich stars, as shown by the black contour. Therefore, binning stars by age in the presence of age uncertainty, results in the disc population appearing noticeably older than it actually is.
@@ Acknowledgements
We thank Sergey Koposov for suggesting the test of the Fourier dipole moment. We thank Keith Hawkins, Sarah Kane, Adam Dillamore and Elliot Davies for their inspirational discussions.

### v3

@@ 6.1.1 The very metal-poor (VMP) disc
Sestito et al. (2019, 2020) found that a large fraction of VMP stars with metallicity $\mathrm{[M/H]}\lesssim-2.5$ have disc-like kinematics. Several possible scenarios for the existence of VMP stars with disc-like orbits are discussed in Sestito et al. (2020). Of these, albeit the least likely and clearly disfavoured by the authors, is that in which the Milky Way disc was already in place during the very metal-poor era. Di Matteo et al. (2020) argued that stars with disc-like kinematics can be found ubiquitously across the entire metallicity range reaching $\mathrm{[M/H]}\sim-6$, thus favouring a scenario of disc formation at very low metallicities. Note that, according to numerical simulations, the MW managed to form a coherently rotating, dominant stellar disc one of the earliest amongst the galaxies of similar mass (see e.g. Semenov et al., 2024; Dillamore et al., 2024a). However, given the mass and metallicity evolution of the Galaxy, its ancient disc stars were not metal-poor.
@@
Figure 10: The effects of the age uncertainty in determining the disc formation time. The left sets of panels are for Au18, and right sets are for Au26. In each sets of figures, the top row focuses on stars on the disc, which $125<R<$ 12 and $|z|<5$ kpc, and the $v_{R}$ vs. $v_{\phi}$ plane is used to distinguish disc and halo population; the bottom row focuses on inner galaxy stars with $R<5$ kpc, and the radial velocity map is used to show the disc star appearance. Left column: the age histograms for the old stars selected from uncertainty scattered age. The distribution of the scattered age is in red, and the true age is in black. The red histogram is truncated because we only focus on the stars we binned with the uncertainty-added age, so the truncation of the red histogram corresponds to the age bin. Middle column: the kinematics of the true old stars born before the spin-up. Right column: the kinematics of the old stars selected with age after scattering with $10\%$ uncertainty. The aqua contours are for highlighting the distribution of all stars in the $v_{R}$ vs. $v_{\phi}$ space, and the black contours are used to illustrate the distribution of metal-rich stars.
@@ 6.1.2 Age of the disc
Queiroz et al. (2023) applied the t-distributed stochastic neighbour embedding (t-SNE) to the chemical abundances measured by APOGEE and used a clustering algorithm to select a population chemically similar to the thick disc. The isochrone ages of the thick disc stars are generally old (their distribution peaks at $\mathrm{age}\gtrsim 10.9$ Gyr). Although their results are not in direct contradiction with our constraints on the disc formation time, we discuss and demonstrate in Section 6.2 below that age uncertainty can produce an overestimation of the disc age measurements. Similarly, Miglio et al. (2021) found an age of $\sim 11$ Gyr after deconvolving the age uncertainty for the chemically-defined thick disc populations with ages from asteroseismological measurements using Kepler data. Gallart et al. (2024) designed a colour-magnitude diagram(CMD)-fitting technique to reveal the star formation history of thin disc stars in the solar neighbourhood ($d\lesssim 100$ pc). They found that the star formation of the disc started around $\sim 11-10.5$ Gyr ago. Belokurov & Kravtsov (2024) separated the globular clusters into in-situ and accreted with an energy-angular momentum cut. The in-situ globular clusters spun up, i.e. increased dramatically their bulk azimuthal velocity at $-1.3\lesssim\mathrm{[Fe/H]}\lesssim-1$ around $11.7-12.7$ Gyr ago, which is also in broad agreement with the result in this work.
Nepal et al. (2024a) used the StarHorse Bayesian isochrone ages (Queiroz et al., 2018) with stellar parameters from the hybrid-CNN analysis of the Gaia RVS sample (Guiglion et al., 2024) and found a kinematically cold, thin disc-like population emerging within 1 Gyr after the Big Bang. Their old, thin disc-like population covers a large range in metallicity from metal-poor to super-metal-rich and is particularly conspicuous at super-solar metallicity. It is difficult to reach super-metal-rich metallicities in the Galactic disc region within 1 Gyr after the Big Bang. It could be therefore that some (or many) of these old disc stars have come from the inner Galaxy, and thus are related to effects of the bar (Dillamore et al., 2023; Yuan et al., 2023; Li et al., 2023; Nepal et al., 2024b; Dillamore et al., 2024b).
We emphasise that our measurements are not influenced by the bar formation time in the Milky Way. In all the Auriga galaxies we inspected, we found that variation in time difference between bar formation and disc formation does not affect our central argument that the oldest stellar population showing bar signatures is the first population born after disc formation. As mentioned in Section 4, the disc population born before bar formation becomes trapped by the Galactic bar, regardless of when the bar formed. Therefore, bar chrono-kinematics are not suitable for distinguishing the bar formation time, whether it happened around $\sim$3 Gyr ago, as suggested by Nepal et al. (2024b), or 7-10 Gyr ago, as suggested by Sanders et al. (2024); Haywood et al. (2024).
@@ 6.2 Effects from the age uncertainty
Scatter due to the age measurement uncertainty causes younger stars to be confused for apparently old stars. As the early Galaxy built its stellar mass up and the star formation rate increased, significantly more stars formed during and after the spin-up of the Milky Way than before. Therefore, the population born before the Galaxy spin-up are a relative minority compared to the subsequent stellar generations. These younger stars can produce noticeable contamination in a given bin of measured old age if the age uncertainty is non-negligible. We demonstrate the effects of such contamination using model galaxies Au18 and Au26 introduced earlier, as shown in Fig. 10.
@@
To illustrate how this contamination can cause overestimation of the disc age, we bin stars with their true ages and the uncertainty scattered ages in an age bin before the spin-up. In the disc region, we use the $v_{R}$ vs. $v_{\phi}$ map to distinguish the halo and the disc population. Focusing on the true age population we see that before the galaxy’s spin-up, no rotation-supported component is found, as expected. However, using the same age bin but now applied to stars with scattered ages, a fast-rotating, disc-like population appears, which can be seen from the aqua contours of the number distribution in the $v_{R}$ vs. $v_{\phi}$ map. This apparently old disc-like component is solely due to the contamination by the younger stars. The disc population is more conspicuous when we further select metal-rich stars, as shown by the black contour. Therefore, binning stars by age in the presence of age uncertainty, results in the disc population appearing noticeably older than it actually is.
@@ Acknowledgements
We thank the reviewer for the helpful comments. We thank Sergey Koposov for suggesting the test of the Fourier dipole moment. We thank Keith Hawkins, Sarah Kane, Adam Dillamore, Elliot Davies, Juntai Shen, and Chengye Cao for their inspirational discussions.